OpenStudy (anonymous):

2x + 7y=17 4x + y= 8

OpenStudy (tad1):

multiply the first equation by -2 to get -4x -14y = -34 then add to the second to get -13y = -26 so y = 2 back substitute to find x 4x _26 = 8 4x =34 x = 17/2

jhonyy9 (jhonyy9):

than y=2 --- 4x +2 = 8 --- 4x = 6 --- x = 6/4 = 3/2 x=3/2 y=2